40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 1 07:30:29 2014, Steve, Update, SUS, ETMX damping restored 
    Reply  Mon Dec 1 16:34:20 2014, Jenne, Update, General, IFO wake-up 
       Reply  Mon Dec 1 17:19:22 2014, Jenne, Update, General, PMC align 
Message ID: 10740     Entry time: Mon Dec 1 16:34:20 2014     In reply to: 10738     Reply to this: 10742
Author: Jenne 
Type: Update 
Category: General 
Subject: IFO wake-up 

 

 After its' several days of rest, it is time to wake up the IFO. 

  • FSS temp was railed at +10, which was making the MC not want to lock.  Set it to zero, locked the MC, and ran the MC WFS relief scripts.
  • PMC trans is down to 0.686, so I'll probably want to tweak that up before I get too carried away for tonight. 
  • c1iscex computer was frozen, which I suspect is why Steve found ETMX tripped this morning.  Soft reboot, and we're back to normal.

With that, it's time for a new week of locking, and trying to catch up with the big kids at the sites.

ELOG V3.1.3-